### v1

2. Armagh Observatory and Planetarium, College Hill, Armagh BT61 9DG, Northern Ireland, UK
Polarimetric and spectropolarimetric observations with FoReRo2: Instrument overview and standard stars monitoring††thanks: Based on data collected with the 2-m RCC telescope at Rozhen National Astronomical Observatory.
In this paper, we present a description and characterisation of the 2-Channel Focal Reducer Rozhen (FoReRo2), with an emphasis on its polarimetric and spectropolarimetric observation modes. FoReRo2 is a multimode instrument mounted at the Ritchey–Chrétien (RC) focus of the 2-meter Ritchey–Chrétien-Coude (RCC) telescope at the Bulgarian National Astronomical Observatory (BNAO) – Rozhen. The primary targets of this instrument include novae, Be/X-ray binaries, symbiotic stars, asteroids, and comets.
Standard stars with a high degree of polarisation are an essential part of the polarimetric observations. In our sample, HD 204827 shows short-term variability in position angle, making it unsuitable for calibration. HD 183143 displays intrinsic polarisation variability but remains stable in position angle. For the first time, we present a statistical study of K and of the Serkowski law, including their mean value, standard deviation, and distribution.
In order to show the FoReRo2 polarimetric capabilities, we also present several examples. The recurrent nova RS Oph shows variability of Serkowski’s K parameter due to dust formation within the first few days after the 2021 outburst. The examples also include imaging polarimetry of comet C/2019 Y4 (ATLAS) and spectropolarimetric observations of the symbiotic star Z And.
Key Words.:
Standards; Techniques: polarimetric; Instrumentation: polarimeters; Stars: binaries: symbiotic; Comets: individual: C/2019 Y4 (ATLAS).1 Introduction
Polarisation is a key characteristic of electromagnetic waves, along with intensity and colour, which opens a new horizon of object characterisation.
The interest in polarimetric and spectropolarimetric observations has increased in recent years, as reflected in several instrumental papers describing various instruments (e.g., WALOP-North (1.3-m telescope, Skinakas Observatory, 2024JATIS..10d4005K), DIPol-UF (Nordic Optical Telescope, 2021AJ....161...20P), CasPol (2.1-5m Jorge Sahade Telescope, Complejo Astronómico El Leoncito, Argentina, 2019JATIS...5b8002S), DIPOL-1 (Sierra Nevada Observatory, 90-cm telescope, 2024AJ....167..137O), SPARC4 (1.6-m telescope, Observatório do Pico dos Dias, 2024BASBr..35...44R), SPHERE/ZIMPOL (VLT, 2018A&A...619A...9S) and ToPol (1.04-m Omicron (West) C2PU facility, Calern Observatory, France) (ToPol)) now replaced by a copy of the DIPol-UF instrument used at NOT (2021AJ....161...20P). This trend is further evidenced by the growing number of observations of standard stars, which play a key role in polarimetric calibration (e.g. 2023A&A...677A.144B; 2024MNRAS.535.1586C; 2024A&A...684A.132M; 2020A&A...635A..46P; 2020BAAA...61B.183L). At higher energies, this renewed emphasis is exemplified by the launch of the Imaging X-ray Polarimetry Explorer (IXPE), which has reopened the X-ray polarimetry window (2016SPIE.9905E..17W).
In this paper, we characterise the polarimetric performances of the 2-Channel-Focal-Reducer Rozhen (hereafter FoReRo2 or FR2 (2000KFNTS...3...13J)). This instrument is attached to the Ritchey–Chrétien (RC) focus of the 2-m telescope of the Bulgarian National Astronomical Observatory (BNAO) – Rozhen. We use a homogeneous dataset of observations of both zero and high polarisation standard stars and present some examples of scientific applications. We assess the stability and the accuracy of the instrument over nearly a decade of observations, and to provide reliable calibration data that can support future scientific use of FR2.
In section 2 we review the FoReRo2 and its observing modes. In section 3, we present the observing strategy and the data reduction steps. Section 4 contains the results of observations of stars with high degrees of polarisation. Section 5 showcases a statistical study of and of polarised stars, spectropolarimetric observations of the recurrent nova RS Oph and the symbiotic star Z And, as well as polarimetric observations (imaging) of the comet C/2019 Y4 (ATLAS). These observations serve as examples to demonstrate FoReRo2’s operational flexibility and its capability to deliver high-quality polarimetric and spectropolarimetric data for a diverse range of astronomical targets.
2 The Two-Channel-Focal-Reducer Rozhen
FoReRo2 was delivered to BNAO based on a contract between the Max-Planck Institute for Solar System Research and the Institute of Astronomy (IA) in 2004. Since that time, the FR2 has been used at the f/8 RC focus of the 2m telescope, where it reduces the original focal ratio to f/2.8, corresponding to an effective focal length of 5.6 m. The FR2 mounted at the telescope is shown in Fig. 1.
The FR2 is a multimode instrument offering a wide variety of observing capabilities:
broadband and narrowband imaging
long-slit, low-dispersion (R) spectroscopy
Fabry-Perot interferometry
imaging polarimetry
low-dispersion spectropolarimetry
A detailed description of the instrument is presented in 2000KFNTS...3...13J. Since that time, several changes and upgrades of the FR2 have been made. Now only two lenses are available, one in the red and one in the blue channel, colour-corrected in the range of 587–1014 nm and 365–436 nm, respectively. A set of Sloan filters and an H filter were added to the set of filters listed in 2000KFNTS...3...13J.
One important improvement of the polarimetric mode is a new Wollaston prism which uses the full aperture of the telescope now. Thus, the previously used quadruple imaging polarimeter, described in 1996Ap&SS.239..259G, was replaced with a dual-beam instrument for polarisation measurements. The Wollaston prism is mounted in a special holder, which has the capability of linear translation (see Fig. 2). This construction allows a quick switch between the different observing modes. In the case of polarimetric and spectropolarimetric observations, the Wollaston prism is inserted into the parallel light beam, and it is taken out of the beam when switching to other observing modes.
The Wollaston Quartz prism is placed in the parallel beam, before the colour divider, thus it feeds both channels simultaneously. The split angle of guarantees separation without overlapping of the ordinary (o) and extraordinary (e) stripes, imaged on the blue and red CCD (see the top panel in Fig. 5).
For several years after the delivery of the FR2 in BNAO, polarimetric measurements were possible in imaging mode only. In 2011, we modified the instrument in such a way that it can also work in spectropolarimetric mode. This was a turning point which opened the way for low-dispersion spectropolarimetry, and the first observation obtained in the spectropolarimetric mode was of comet C/2009 P1 (Garrad)111(http://nao-rozhen.org/observations/fr8.html).
For several years, the polarimetric mode was used by rotating the whole FR2 to different position angles in order to obtain the necessary and sufficient set of observations needed for calculating the Stokes parameters Q and U. In February 2014, an ”Agreement for collaborative research actions between Armagh Observatory (AO) and Institute of Astronomy and National Astronomical Observatory (IA and NAO) at the Bulgarian Academy of Sciences (BAS)” was signed. In the framework of this collaboration, AO delivered to BNAO a /2 retarder plate (half-wave plate (HWP)) for joint polarimetric observations with the 2-m telescope. The design and characteristics of waveplates can be found in 2004JQSRT..88..319S. The waveplate used with FR2 is a Super-Achromatic True Zero-Order Waveplate 5 - APSAW-5. More details are presented on the webpage of the vendor222http://astropribor.com/super-achromatic-quarter-and-half-waveplate/. A dedicated, precise mechanical construction and electronic control system for the rotation of the retarder was designed and manufactured in the workshop of the Institute of Astronomy and BNAO. All components are mounted on a plate, which is shown in Fig. 3. On the bottom side of the plate (not seen in the figure), a gear is mounted that holds the rotation cage (the housing of the HWP has a clear aperture of 40 mm). A worm drive gear, rotated by a stepping motor, is coupled to the cage.
The diameter of the /2 retarder is 40 mm, and it is mounted immediately below the RC focal plane (marked as HWP in Fig 2). A better solution would be to mount the /2 retarder in the parallel beam, but as the diameter of the latter is 50 mm, such a solution would cause light losses greater than 35%.
The rotation of the /2 retarder is controlled remotely with a positional accuracy better than . This accuracy was achieved with precise mechanical construction and control electronics consisting of an Arduino board fitted with a Stepper Motor Driver.
The Arduino board runs the motor control and precise positioning firmware, which is accessed over a serial interface by a Raspberry Pi single-board computer. The Raspberry Pi computer runs the user software to control the angle of the /2 retarder plate and is accessed remotely over the Local Area Network.
In order to avoid any backslashes in the gears, the firmware is designed in such a way that it always approaches the needed position from one direction only. The absolute angle rotation is achieved using a precise optical null position indicator, which is turned on only during the rotation in order to avoid any light spill during the exposure. The null position indicator is also used for self-calibration.
The delivered HWP had no mark showing the orientation of its fast axis. Therefore, we performed a laboratory experiment with a twofold aim: (1) to identify the orientation of the fast axis, and (2) to check the mechanical uniformity over a full cycle of rotation. For this experiment, a white screen, inclined at with respect to the optical path, was inserted above the RC focal plane. It was illuminated with a tungsten lamp, and the reflected light passed through the HWP. After that, the o- and e-beams were separated by the Wollaston prism. The contrast between the o- and e-beams was increased by putting a Glan prism in the RC focal plane, i.e. between the reflecting screen and the HWP. The Glan prism was rotated manually until the best contrast was achieved (at angle of the HWP). After that, the HWP was rotated in steps of , covering the full range of angles from to . The signal of the o- and e- beams, and , respectively, was measured in rectangles of equivalent size 5661 px, positioned around the centres of their respective image stripes. The result of this laboratory experiment is shown in Fig. 4. The three panels contain dependencies of the rotation angle of the fast axis of the HWP. In the top panel of this figure, the measured signals of both beams, and , are shown, as squares and triangles, respectively. The middle panel shows the calculated values of the quantity and its approximation with a function. The bottom panel shows the difference between the measured values and the fit. These residuals exhibit substantial fluctuations in uniformity during the full cycle of rotation. One relatively uniform region, close to the difference (data-fit) of , is seen in the range from to . We should note that this laboratory experiment is used with the FR2 only, not with the entire system telescope plus focal reducer.
In 2018, both channels of the FR2 were equipped with new detectors. These are Andor CCD cameras iKon-L 936333https://andor.oxinst.com/products/ikon-ccd-cameras-for-long-exposure, with the new dual antireflection deep depletion ’BEX2-DD’ chip. One great advantage of the new cameras is the thermo-electric cooling, and another one, especially for the red channel, is the suppression of the fringing in the red and near-IR spectral region. The cameras comprise 20482048 quadratic pixels, each having a size of 13.5 m with a resulting scale of 0.497 arcsec/px and a FoV of about 17x17 arcmin (in imaging mode).
For spectropolarimetric observations, we use a grism with 300 lines in the red channel. Two slit widths are available: 110 m and 200 m. Using the 110 m slit yields a resolving power of at , with a reciprocal spectral dispersion of 3 pixel-1. Over the past eight years (after recoating the 2 m mirror) in order to obtain more accurate results, we used only slitless spectropolarimetry (See Section 3.3.1 for details). In the slitless mode, the resolving power depends on the seeing.
Raw images of the polarimetric (top: comet C/2019 Y4 (ATLAS)) and spectropolarimetric (bottom: Be/X-ray binary star X Per) mode of observations are presented in Fig. 5. For the polarimetric mode of observations, a mask of size 35 mm 4 mm is used. For slitless spectropolarimetric observation, a mask of size 4 mm 3 mm is used. Two spectra of Be/X-ray binary star X Per are presented in Fig. 5, where the emission line is visible. The spectra correspond to the o- and e- beam, respectively.
3 Observing strategy and data reduction
3.1 Observations and basic reduction steps
We obtain polarised spectra and polarised images at eight HWP angles apart.
Standard stars with zero degrees of polarisation were used to correct the instrumental polarisation. Standard stars with a high degree of polarisation were used to correct the position angle. Polarisation standards were observed with the same instrumental setup as astronomical objects on the same nights.
The spectra are reduced using IRAF (1993ASPC...52..173T) in the standard way, including bias removal and wavelength calibration. For the spectropolarimetric data reduction, our research group created dedicated IRAF scripts. For better comparison with catalogue values in the V filter, we used the extracted and from spectra using the IRAF SBAND procedure.
The process of spectropolarimetric data reduction includes the following: bias subtraction; extracting 1D spectra; wavelength calibration; beam swapping technique; corrections for instrumental polarisation, chromatism of the HWP, and position angle correction. The beam swapping technique (BST) (2009PASP..121..993B) is the core of the sequence of reduction steps. We use it to minimise the instrumental polarisation.
We tested spectropolarimetric data reduction with and without the use of flat fields. Our tests show that in both cases, with and without flat fields, the results obtained are indistinguishable within the precision of our data. This result corresponds to the result obtained by other authors (2019JATIS...5b8002S).
3.2 Beam swapping technique
A beam-swapping technique is used for polarimetric data processing. The values of F and , where and F are the fluxes of the o- and e- beams, respectively, are obtained for the different HWP angles. We define , where corresponds to the HWP angle.
The following formulae are used to calculate the Stokes parameters (2009PASP..121..993B):
Figure 6 presents the Stokes Q and U parameters of the unpolarised standard star HD 154892 after applying the beam-swapping technique. As Fig. 6 demonstrates, after applying the BST, the observed polarisation of the unpolarised standard star is close to zero.
3.3 Instrumental polarisation
Regular spectropolarimetric observations started after the installation of the HWP in November 2015. The BST can reduce instrumental polarisation caused by the optical components located between the HWP and the CCD cameras. The residual instrumental polarisation is introduced by the optical elements before the HWP and the combination of the reflective coating of the primary and secondary mirrors. This residual instrumental polarisation was corrected using standard stars with zero degree of polarisation. We correct the Stokes parameters and of the objects for instrumental polarisation using the Stokes parameters and of standard stars with zero degree of polarisation using the following equations:
where and represent both Stokes parameters corrected for the instrumental polarisation.
The 2m mirror’s aluminium reflective coating has to be renewed from time to time. The last two aluminium reflective coatings occurred in 2008 and 2017. Figure 7 shows the Q-U diagram of zero polarisation standard stars in a synthetic V filter before and after recoating of the 2m mirror. A synthetic V filter represents the standard Johnson V-band transmission curve and is used in synthetic photometry to simulate photometric measurements from spectral data. When applied (e.g., via IRAF’s SBAND task), it multiplies the spectrum by the filter transmission curve and integrates the result to estimate the flux or magnitude as if observed through a real V-band filter. This allows a direct comparison between the spectro-polarimetric and polarimetric data.
In Figure 7, different symbols indicate the spectropolarimetric observation modes (with and without slit, before and after the recoating of the 2-m mirror), while the colour represents the modified Julian date (MJD). When a slit is used (triangles and squares), the instrumental polarisation is high and, before the mirror recoating, also highly variable (triangles). In the slitless mode, no significant difference is observed before and after the recoating. The mean Stokes parameters in the synthetic V filter for the slitless mode are and , marked with the black pentagon. The ellipses represent the 1-, 2-, and 3- standard deviations. The diagram also shows a shift of the slitless-mode points (a decrease in Stokes Q), with the latest data lying to the left of the black pentagon, possibly due to changes in instrumental polarisation and the selection of more accurate zero-polarisation standard stars.
Figure 8 shows the polarised spectra of the unpolarised standard star HD 191195 over the wavelength range 4800–8200 Å. No wavelength dependence is observed in the Stokes parameters. The widths of the orange and blue bands indicate the standard deviation from the mean values, calculated over the entire wavelength range. Multiple reflections within the wave plate (e.g., 2003SPIE.4843..437I; 2006AJ....132..433K) can produce the so-called ripple effect; however, no significant (¿0.1%) ripple pattern is detected in the observed Stokes Q and U.
3.3.1 Slit vs. slitless spectropolarimetry
Two types of observations of zero and high polarisation standards were performed - slit and slitless spectropolarimetric observations before and after recoating of the 2m mirror, respectively.
Results of slit- and slitless spectropolarimetric observations before and after recoating of the 2m mirror are presented in Fig. 9 and Fig. 10, respectively. Before the recoating of the 2m mirror, results of slit- and slitless spectropolarimetric observations are practically identical (Fig. 9). After the recoating of the 2m mirror results of slit- and slitless spectropolarimetry observations are different (Fig. 10).
It is known that when the width of a slit is comparable to the wavelength of the incident light, the slit acts as a partial polariser. Even though in our case the slit width is around 150 times the Hα wavelength by using the theory by Slater42, based on microwave wave-guide theory, which also includes the depth of the slit and its material conductivity, we can show that the instrumental polarisation caused only by the slit is in the order of a few percent. This is well explained by AstroSpol-Keller in the book “AstroSpol” ed. by AstroSpol.
In our case, this addition to the instrumental polarisation is most likely due to the combination of the explained slit effect, as well as a small misalignment of the 2m mirror relative to the optical axis, which leads to a slightly deformed star PSF (coma), as well as the narrow metal slit, which cuts the star PSF asymmetrically. This is exactly what you can see on Figures 7 and 10. Before and after the mirror re-aluminization, the Stokes Q has a different value, which is in the order of 1%. Figure 10 even shows how the Stokes U is affected by the instrumental polarisation caused by the coma-slit combination and its wavelength dependence.
Results of slitless spectropolarimetric observations are presented in Section 4. Only when using slitless spectropolarimetric observations we could properly correct for the instrumental polarisation.
3.4 Correction for the chromatism (wavelength dependence) of the half-wave plate
For correction of the chromatism of the HWP, we used the wavelength dependence function , which is represented in Fig. 11(top panel). This function was derived from the wavelength dependence of the observed position angle of the high polarisation standard star HD 161056. The achromatic function is well described by a sixth-degree polynomial fit:
We used the following equations to correct for the chromatism (wavelength dependence) of the HWP.
where and represents Stokes parameter corrected for the chromatism. We used this values of Stokes parameters to calculate . The results of this correction are represented in Fig. 11(bottom panel) where gray data represent the position angle before chromatism correction, while the blue data show it after correction.
3.5 Correction for position angle.
The final correction applied during data reduction is the correction for the position angle. Standard stars with a high degree of polarisation were used to correct the position angle. represents the difference between the observed values in the synthetic V filter and the catalogue value. The correction follows the equation:
where represents the difference between the observed and catalogue values of the position angle. The average value of . The RMS of is , which indicates the instrument’s stability with respect to the position angle correction. In figure 12 we present the observed degree of polarisation in the synthetic V band and . The vertical columns represent catalogue values of the degree of polarisation in the V band, and the width of the column corresponds to the uncertainties in the catalogue values. The horizontal grey line represents the average value of . This offset of P.A. depends on the instrument alignment as well as asymmetric ageing of the mirror coating from season to season. For this reason, it is necessary to determine these P.A. offsets for each observation night or each observation season.
3.6 Selection of aperture for imaging polarimetry.
Imaging aperture polarimetry was performed as explained in 2011JQSRT.112.2059B by selecting the aperture at which the reduced Stokes parameters and converge to a well-defined value.
This method differs from the standard aperture photometry with apertures of increasing size in the way we chose the best aperture size. Here, we are looking at which aperture size the ratios between the fluxes in the parallel and the perpendicular beams (hence the measured Stokes parameters) become constant. This aperture is usually smaller than the values where the individual fluxes become asymptotic. Technically, the best aperture is selected using plots like those of Figure 13
The results of the imaging polarimetry measurements through the years of observations with the polarimetric mode of the FoReRo2 instrument are compared with their catalogue values and presented in Figure 14 and Appendix A within Tables 5 and 6 therein.
4 Spectropolarimetric observations of high polarisation standards
In the spectropolarimetric mode of observation, we observed four unpolarised standard stars: HD 191195, HD 21447, HD 154892, and HD 212311. We also observed ten highly polarised standard stars: HD 7927, HD 25443, HD 161056, HD 19820, HD 204827, BD +25 727, HD 183143, BD +59 389, HD 154445, and HD 155197. The catalogue and observed values of these stars are presented in Table 1 and their polarised spectra are presented in Figure 18. The position angle of the highly polarised standard stars is used as a reference. In Fig. 19, we present 3D maps of these stars. The degree of polarisation is indicated by the length of the orange lines, while the orientation of each line represents the position angle. The colour of each star corresponds to its distance.
Table 1 includes the following information: the observed unpolarised and polarised standard stars, their coordinates, the number of observations for each star, the average value in the synthetic V band, the standard deviation, the degree of polarisation and position angle from catalogues, and references. Two highly polarised standard stars, HD 183143 and HD 204827, exhibited variability during our observations.
| Name | R.A. | Dec. | № of | 1 | 2 | Band | P.D. (%) | P.A. (deg.) | References |
| (J2000) | (J2000) | obs. | (%) | (%) | (catalogue) | (catalogue) | |||
| Unpolarised standard stars | |||||||||
| HD 191195 | 20 06 13.9 | +53 09 56.5 | 10 | 0.03 | 0.02 | – | 0.004 0.001 | (b) | |
| HD 21447 | 03 30 00.2 | +55 27 06.5 | 16 | 0.05 | 0.02 | V | 0.051 0.020 | (a) | |
| HD 154892 | 17 07 41.3 | +15 12 37.6 | 20 | 0.08 | 0.04 | B | 0.050 0.030 | (c) | |
| HD 212311 | 22 21 58.6 | +56 31 52.7 | 10 | 0.11 | 0.03 | V | 0.034 0.021 | (a) | |
| High polarisation standards | |||||||||
| HD 7927 | 01 20 04.9 | +58 13 53.8 | 3 | 3.25 | 0.08 | V | 3.298 0.025 | 91.08 0.20 | (a) |
| HD 25443 | 04 06 08.1 | +62 06 06.6 | 14 | 5.21 | 0.05 | V | 5.127 0.061 | 134.23 0.34 | (a) |
| HD 161056 | 17 43 47.0 | -07 04 46.6 | 27 | 4.05 | 0.04 | V | 4.030 0.025 | 66.93 0.18 | (a) |
| HD 19820 | 03 14 05.3 | +59 33 48.5 | 11 | 4.82 | 0.03 | V | 4.787 0.028 | 114.93 0.17 | (a) |
| BD +25 727 | 04 44 24.9 | +25 31 42.4 | 2 | 6.31 | – | V | 6.280 0.003 | 32.20 0.10 | (f) |
| BD +59 389 | 02 02 42.1 | +60 15 26.4 | 2 | 6.63 | – | V | 6.701 0.015 | 98.09 0.07 | (a) |
| HD 154445 | 17 05 32.3 | -00 53 31.4 | 3 | 3.72 | 0.06 | V | 3.780 0.062 | 88.79 0.47 | (a) |
| HD 155197 | 17 10 15.8 | -04 50 03.7 | 1 | 4.45 | – | V | 4.320 0.023 | 102.84 0.15 | (a) |
| Variable high polarisation standards | |||||||||
| HD 183143 | 19 27 26.6 | +18 17 45.2 | 7 | 6.13 | 0.15 | V | 6.100 0.050 | 179.30 0.20 | (d) |
| HD 204827 | 21 28 57.8 | +58 44 23.2 | 9 | 5.56 | 0.20 | V | 5.322 0.014 | 58.73 0.08 | (a) |
| V | 5.607 0.003 | 58.33 0.02 | (e) | ||||||
References: (a) 1992AJ....104.1563S; (b) 2020A&A...635A..46P; (c) 1990AJ.....99.1243T;
(d) 1982ApJ...262..732H;
(e) 2021AJ....161...20P; (f) Nordic Optical Telescope3
1
The average degree of polarisation in the synthetic V band
2
The standard deviation calculated as in Eq. 8
- 3
The average polarisation deviation between the catalogue values and the mean values measured for high-polarisation standard stars in the synthetic V band is (. This value is comparable to the standard deviation for the same set of stars, as listed in Table 1. Therefore, can be interpreted as a representative measure of the instrument’s polarimetric accuracy in the synthetic V band. The stars HD 183143 and HD 204827 were excluded from this calculation due to their known polarimetric variability.
4.1 Searching for variability.
Usually, the standard deviation is used to identify small fluctuations in magnitude and detect flickering (e.g., 2023BlgAJ..38...83Z). In this paper, we apply the same approach to detect and confirm the variability of highly polarised standard stars.
The standard deviation is calculated as:
where is the average degree of polarisation in the synthetic V band, is the number of data points. The calculated in this way includes the variability of the degree of polarisation of the star (if it exists).
In Table 1, the standard deviation is calculated for stars with more than two observations. The typical error of spectropolarimetric observations (presented in Figure 18) is approximately 0.05–0.1%. The standard deviations are comparable to this value. The standard deviation of unpolarised standard stars serves as an indicator of instrumental stability, while the standard deviation of highly polarised standard stars indicates the accuracy of the methodology. Two of the observed highly polarised standard stars, HD 183143 and HD 204827, demonstrate a higher standard deviation: 0.15% and 0.20%, respectively. Both stars exhibit variable polarisation, which indicates the presence of an intrinsic component of polarisation.
Spectropolarimetric observations of HD 183143 with FoReRo2 span a period of two years, from 4 June 2022 to 24 October 2024. The star exhibits an emission line and shows a variable degree of polarisation, indicative of an intrinsic polarisation component. The amplitude of variability in the synthetic V band is . On each of the seven observing nights, another highly polarised standard star was observed alongside HD 183143. The position angle of HD 183143 was corrected using the reference value for the corresponding night. Despite the variation in polarisation degree, the position angle remains stable, with a mean value in the synthetic V band of . Therefore, HD 183143 may still be used as a reference for position angle calibration. It is worth noting that the position angle we derive differs by approximately from the catalogue value, which may indicate long-term variability in the polarisation angle.
1986PASP...98..792D concluded that HD 204827 possesses a component of intrinsic polarisation. Our observations confirm the variability of this star. HD 204827 was observed on two consecutive nights, 2 and 3 May 2025, with position angle calibration performed using a second high-polarisation standard on each night (HD 161056). The applied position angle corrections were and for 2 and 3 May, respectively. HD 204827 exhibits variability in both the degree of polarisation and the position angle. In the synthetic V band, we obtain , on 2 May 2025, and , on 3 May 2025. Despite careful calibration, the derived position angle of HD 204827 differed by nearly between the two nights, with no identifiable instrumental cause. This observed variability suggests that HD 204827 is not a reliable standard for polarimetric calibration, particularly with respect to position angle stability.
5 FoReRo2 observational polarimetric examples
As illustrative examples, we show a statistical study of and based on the spectropolarimetric observations of polarised stars, spectropolarimetric observations of the recurrent nova RS Oph and the symbiotic star Z And, and the imaging polarimetry of the comet C/2019 Y4 (ATLAS).
5.1 Serkowski law: relationship between and
5.1.1 Fitting the polarisation spectra with the Serkowski law
The Serkowski law is empirical and describes the wavelength dependence of linear polarisation (1975ApJ...196..261S):
where is the peak of the interstellar polarisation at wavelength . The K coefficient describes the shape of the curve, and in Serkowski’s original work (Serkowski et al. 1975), it has a value of 1.15. The parameters of Serkowski’s law fit with catalogue values, and values from this work are presented in Table LABEL:tab.Serkowski. In Fig. 18, polarisation spectra of these high degree of polarisation standards are presented.
| High polarisation standards | (%) | () | ||
| HD 7927 | 3.25 | |||
| HD 25443 | 5.21 | |||
| HD 161056 | 4.05 | |||
| HD 19820 | 4.82 | |||
| BD +25 727 | 6.31 | |||
| BD +59 389 | 6.63 | |||
| HD 154445 | 3.72 | |||
| HD 155197 | 4.45 | |||
| Variable polarisation standards | (%) | K | () | |
| HD 183143 | ||||
| 2023-08-13 | 6.3 | |||
| 2024-04-08 | 5.92 | |||
| HD 204827 | ||||
| 2019-07-02 | 5.67 | |||
| 2019-07-01 | 5.23 |
5.1.2 A statistical study of and : mean, standard deviation, and data distribution.
Initially, K = 1.15 was assumed as a standard 1973IAUS...52..145S. An attempt to find a relationship between the parameters of Serkowski law, and can be found in the works of 1982AJ.....87..695W; 1992ApJ...386..562W. In this section, we apply a statistical analysis of and to determine the mean, standard deviation, and data distribution.
1992ApJ...386..562W suggested the relationship:
where is expressed in m. 2017MNRAS.464.4146C, when comparing the K and values presented in his paper with those reported by 1992ApJ...386..562W, reveals a systematic and statistically significant deviation, with the K values reported by 1992ApJ...386..562W, most likely being shifted due to incorrect effective wavelengths for the R and I passbands (sec. 4.2.1. in 2017MNRAS.464.4146C).
Since spectropolarimetric data do not suffer from the problem of properly characterising the photometric system (2017MNRAS.464.4146C), we decided to use only spectropolarimetric data published in the literature, together with data obtained with FR2, to perform a statistical analysis of the values in the plane. We used data published in (2017A&A...608A.146B; 2017MNRAS.464.4146C; 2018A&A...615A..42C; 2022NewA...9701859N; 2023eas..conf.1422N). The coefficients and for nonvariable standard stars are presented in Table LABEL:tab.Serkowski to calculate the mean and standard deviation. For calculations, we used 146 stars, 31 of which were obtained with FR2. The mean of is with a standard deviation of m. The mean of the K coefficient is 1.23 with a standard deviation of 0.32. It is noteworthy that the mean value of K is slightly higher than 1.15, the initial value published by 1973IAUS...52..145S.
In Fig. 15 we present values of Serkowski’s fit of spectropolarimetric data. We present the variable and of RS Oph (with purple) and HD 204827 (with dark blue). With black is presented the sample of Serkowski fit to SNe Ia, taken by (2017ApJ...836...88Z) and Appendix B in (2018A&A...615A..42C). We do not use these values to calculate the mean and standard deviation. In the plane, the spectropolarimetric standard stars (obtained in this work and 2017MNRAS.464.4146C) are located upper to the relationship determined in 1992ApJ...386..562W. Our observations of the standard stars (with red in Figure 15) show the same systematic discrepancy. The orange ellipses present the areas with ; , and . Most of the data is located in . Two of the values for the K coefficient of RS Oph, 2 and 6 days after the most recent outburst, are located outside of . The biggest deviation is observed in the sample of SN Ia (2017ApJ...836...88Z).
5.2 Variable and of the recurrent nova RS Oph.
Spectropolarimetric observations of the recurrent nova RS Oph following its latest outburst indicate a significant intrinsic component of polarisation (2023A&A...679A.150N). In Table 3, we present the fit of the spectropolarimetric observations of RS Oph using Serkowski’s law. The peak value of the K coefficient is 1.80. From day 2 to day 7 after the 2021 outburst of RS Oph, the K coefficient decreases. It is noteworthy that the peak of the K coefficient within the first week of the most recent outburst of RS Oph corresponds to the newly formed dust during the event. This dust was destroyed within nine days of the 2021 outburst (2023A&A...679A.150N). In Table 3 we present the data with a correlation coefficient higher than 0.75. The variable K coefficient of RS Oph is presented in purple in Figure 15.
The variation in RS Oph is dominated by changes in the Serkowski K coefficient, while remains essentially constant. This suggests that the intrinsic polarisation, due to nova-related dust, affects the shape of the polarisation curve without altering the peak wavelength. In contrast, some supernovae exhibit both high K values and unusually short (e.g., (2017ApJ...836...88Z)), indicating either a different dust grain population or a more complex combination of intrinsic and interstellar polarisation components. Therefore, the behaviour of RS Oph cannot fully account for the polarimetric properties seen in supernovae, but it provides a reference case for transient dust-induced intrinsic polarisation.
| Date | Days into the | K | Correlation | |||
| 2021 the outburst | m | coefficient | ||||
| 2019-07-02 | -768 | 0.563 0.001 | 2.90 0.01 | 1.21 0.01 | 0.79 | |
| 2021-08-10 | 2 | 0.564 0.000 | 3.09 0.01 | 1.80 0.01 | 0.96 | |
| 2021-08-14 | 6 | 0.575 0.000 | 3.10 0.01 | 1.77 0.01 | 0.92 | |
| 2021-08-15 | 7 | 0.567 0.000 | 3.08 0.01 | 1.49 0.01 | 0.92 | |
| 2022-06-04 | 300 | 0.554 0.001 | 2.93 0.01 | 1.45 0.01 | 0.78 | |
| 2024-04-08 | 974 | 0.574 0.000 | 2.88 0.01 | 1.49 0.01 | 0.84 | |
| 2024-05-05 | 1001 | 0.536 0.000 | 3.11 0.01 | 1.43 0.01 | 0.92 |
5.3 Spectropolarimetric observation of symbiotic star Z And
The spectropolarimetric observations of the two Raman OVI lines at and are powerful tools to diagnose the geometry of the symbiotic stars (1997A&A...321..791S; 1996A&A...310..235H; 1997A&A...327..219S). These emission lines are due to Raman scattering of the OVI resonance doublet at and by neutral hydrogen (1989A&A...211L..31S). Narrow-band Raman OVI imaging is used as a photometric tool for hunting symbiotic stars in the local Universe (2019AJ....157..156A).
Based on spectropolarimetric observations of the Raman lines, 1997A&A...327..219S determined the orbit inclination and the orbit orientation of Z And. We use this symbiotic star as a sample to explore the possibilities of FoReRo2 in the spectropolarimetric mode of observations. The symbiotic star Z And was observed on August 22, 2020 and August 10, 2021. Fig. 16 shows normalised intensity (Stokes I) and observed degree of polarisation (top panel); null parameter (middle panel); and observed position angle (bottom panel). The polarisation at and is well visible.
The null parameters are described in 2009PASP..121..993B. Both parameters are used as a quality check tool for data reduction. In Fig. 16 (middle panel) both parameters: and are close to zero at . That means the observed degree of polarisation is real.
For the wavelength calibration, one may use Ne, HgCd, and He lamps. Our experience with the spectropolarimetric data reduction shows that wavelength calibration must be applied after splitting the spectra of o- and e-beam (both for object and arc lamp) into separate fits files. Otherwise, small wavelength shifts between o- and e-beam lead to a dummy degree of polarisation in the spectral lines (for more information about small wavelength shifts and their impact on line polarisation see 2013A&A...559A.103B).
5.4 Imaging polarimetry of comet C/2019 Y4 (ATLAS)
Imaging polarimetry of comet C/2019 Y4 (ATLAS) in the wide-band filter Rx (=694.0 nm, FWHM=79.0 nm), which is chosen in such a way that the molecular emission to be avoided and only the reflected sunlight from the dust to pass through, was carried out during 7 nights in April and May 2020 in a wide range of phase angles above 55∘ during its disintegration. To select the aperture for imaging polarimetry, we use the methodology described in Subsection 3.6.
The measured degree of linear polarisation for each night is presented in Table 4. These results were fitted with the function from Eq. 11 and presented together with measurements from the comets polarimetry database (2010PDSS.8124E....K) in Figure 17.
The obtained coefficients from the fit are as follows:
Nevertheless, the disintegration of its nucleus, comet ATLAS, shows polarisation behaviour similar to any ordinary comet. Its phase-polarisation curve has a maximum linear polarisation LPmax=21.9 % at phase angle of 88.6∘.
The slope at the inversion angle is computed by Eq. 12 and is h= %/deg.
| Date444Date of observations, DD/MM/YYYY | S-T-O555Phase angle (Sun-Target-Observer) in degrees | LP666Linear polarisation in % | 7771- uncertainty of LP in % |
| 18/04/2020 | 55.90 | 12.331 | 0.018 |
| 25/04/2020 | 62.25 | 12.488 | 0.069 |
| 27/04/2020 | 64.25 | 12.812 | 0.058 |
| 30/04/2020 | 67.50 | 14.340 | 0.069 |
| 05/05/2020 | 73.80 | 16.901 | 0.068 |
| 12/05/2020 | 86.20 | 21.965 | 0.041 |
| 13/05/2020 | 88.50 | 22.119 | 0.060 |
6 Summary and conclusions
In this study, we present an overview of the 2-Channel Focal Reducer Rozhen (FoReRo2), with emphasis on its polarimetric and spectropolarimetric observing modes. To assess the instrument’s performance and stability, we analysed spectropolarimetric observations of four unpolarised standard stars over 62 epochs, and ten high-polarisation standard stars over 79 epochs. The dataset spans a decade of polarimetric and spectropolarimetric observations with FoReRo2.
The typical polarimetric accuracy of FoReRo2, estimated from high-polarisation standard stars, is better than 0.1% in the synthetic V band as well as in the red domain from imaging polarimetry.
The standard deviation on long-term observations of high polarisation standards can be used as a tool to detect the variability of the degree of polarisation. We used this methodology to demonstrate the variability of two high polarisation standard stars: HD 183143 and HD 204827.
Among the ten high-polarisation standard stars analysed in this work, HD 204827 exhibited significant short-term variability in its position angle. Based on our observations, we conclude that HD 204827 is unsuitable for use as a reliable polarimetric standard.
Based on spectropolarimetric observations, we made a statistical analysis of the Serkowski law parameters ( and ). The mean of at 0.58 with a standard deviation of . The mean of the coefficient is 1.23 with a standard deviation of 0.32.
Selection of the aperture for imaging polarimetry is performed by choosing the aperture at which the reduced Stokes parameters ( and ) converge to a well-defined value.
Only slitless spectropolarimetry gives reasonable results, while normal slit spectropolarimetry seems affected by strong instrumental polarisation. This is likely due to a slight misalignment of the 2-meter mirror relative to the optical axis appears when it was mounted after its last recoating.
The 2021 outburst of RS Oph demonstrates that intrinsic polarisation caused by nova-related dust can modify the Serkowski K parameter, with no significant change in . This provides a reference case for dust-induced polarimetric variability in transient events.
The symbiotic star Z And demonstrates variable polarisation in the Raman OVI line at .
The disintegrated comet ATLAS, shows maximum linear polarisation LPmax=21.9 % at phase angle of 88.6∘, similar to any ordinary comet.
To our knowledge, FoReRo2 is the only operational low-resolution optical spectropolarimeter mounted on a 2-meter-class telescope in Southeastern Europe. This work presents the first comprehensive calibration of the instrument and demonstrates its long-term stability and scientific applicability. The results provide a reliable reference baseline for future polarimetric and spectropolarimetric observations with FoReRo2. The instrument is well suited for transient phenomena, as well as for regular studies of solar system bodies, the interstellar medium, and various stellar sources.
Acknowledgements.
Y.N. acknowledges partial support by grants: K-06-H58/3 ”Space weather and habitability of exoplantes: the role of coronal mass ejections and (super)flares” and K-06-M88/1 ”Interstellar medium - 3D distribution and physical characteristics”, by the Bulgarian National Science Fund. G.B., Y.N. and T.B. acknowledge partial support by grants: K-06-H88/5 “Physical properties and chemical composition of asteroids and comets - a key to increasing our knowledge of the Solar System origin and evolution.” by the Bulgarian National Science Fund. Yanko Nikolov, thanks for the hospitality of Armagh Observatory and Planetarium, and for the opportunity to work there for a period of several weeks. The research that led to these results was partially carried out with the help of infrastructure purchased under the National Roadmap for Research Infrastructure, financially coordinated by the Ministry of Education and Science of Republic of Bulgaria. Astrometry.net was used to derive the scale of the system with the new CCD cameras. This research has made use of NASA’s Astrophysics Data System Bibliographic Services.Appendix A Imaging polarimetry of standard stars
This appendix presents imaging polarimetry results for both unpolarised and polarised standard stars. Table 5 summarises the measurements for unpolarised standards, while Table 6 lists the results for polarised standard stars. The data illustrate the instrument’s performance in imaging polarimetric mode across various bands. Some of the observations were conducted using narrowband filters.
| Star | Date | Filter | P, | P, | Pcat, | Pcat, | (P-Pcat), | (P-Pcat), | Notes |
| HD42807 | 2014-12-20 | V | 0.07 | 0.01 | 0.00 | 0.00 | 0.07 | 0.01 | UKIRT; 000067 |
| G191B2B | 2017-12-15 | R | 0.13 | 0.05 | 0.09 | 0.05 | 0.04 | 0.07 | NOT; 000067 |
| G191B2B | 2017-12-15 | R | 0.20 | 0.09 | 0.09 | 0.05 | 0.11 | 0.10 | NOT; 090157 |
| HD154892 | 2018-03-11 | R | 0.07 | 0.00 | 0.05 | 0.03 | 0.02 | 0.03 | NOT; 000067 |
| HD154892 | 2018-03-11 | R | 0.12 | 0.00 | 0.05 | 0.03 | 0.07 | 0.03 | NOT; 090157 |
| – | 2018-10-02 | R | 0.04 | 0.03 | 0.00 | 0.00 | 0.04 | 0.03 | 000067 |
| G191B2B | 2019-09-21 | Rx | 0.13 | 0.05 | 0.09 | 0.05 | 0.04 | 0.07 | NOT; 000067 |
| G191B2B | 2019-09-21 | Rx | 0.09 | 0.05 | 0.09 | 0.05 | -0.01 | 0.07 | NOT; 090157 |
| G191B2B | 2019-10-21 | Rx | 0.11 | 0.05 | 0.09 | 0.05 | 0.02 | 0.07 | NOT; 000067 |
| G191B2B | 2019-10-21 | Rx | 0.20 | 0.05 | 0.09 | 0.05 | 0.11 | 0.07 | NOT; 090157 |
| G191B2B | 2019-10-22 | Rx | 0.03 | 0.05 | 0.09 | 0.05 | -0.06 | 0.07 | NOT; 000067 |
| G191B2B | 2019-10-22 | Rx | 0.09 | 0.05 | 0.09 | 0.05 | -0.01 | 0.07 | NOT; 090157 |
| G191B2B | 2019-10-24 | Rx | 0.11 | 0.06 | 0.09 | 0.05 | 0.02 | 0.08 | NOT; 000067 |
| G191B2B | 2019-10-24 | Rx | 0.06 | 0.06 | 0.09 | 0.05 | -0.03 | 0.08 | NOT; 090157 |
| G191B2B | 2019-10-25 | Rx | 0.09 | 0.07 | 0.09 | 0.05 | -0.00 | 0.09 | NOT; 000067 |
| G191B2B | 2019-10-25 | Rx | 0.06 | 0.08 | 0.09 | 0.05 | -0.03 | 0.09 | NOT; 090157 |
| G191B2B | 2020-04-18 | Rx | 0.10 | 0.06 | 0.09 | 0.05 | 0.01 | 0.07 | NOT; 000067 |
| G191B2B | 2020-04-18 | Rx | 0.12 | 0.06 | 0.09 | 0.05 | 0.03 | 0.07 | NOT; 090157 |
| G191B2B | 2020-04-25 | Rx | 0.05 | 0.05 | 0.09 | 0.05 | -0.04 | 0.07 | NOT; 000067 |
| G191B2B | 2020-04-25 | Rx | 0.04 | 0.05 | 0.09 | 0.05 | -0.06 | 0.07 | NOT; 090157 |
| HD154892 | 2020-04-30 | Rx | 0.08 | 0.04 | 0.05 | 0.03 | 0.03 | 0.05 | NOT; 000067 |
| HD154892 | 2020-04-30 | Rx | 0.05 | 0.04 | 0.05 | 0.03 | 0.00 | 0.05 | NOT; 090157 |
| HD154892 | 2020-05-01 | Rx | 0.06 | 0.03 | 0.05 | 0.03 | 0.01 | 0.04 | NOT; 000157 |
| – | 2022-06-04 | R | 0.06 | 0.04 | 0.00 | 0.00 | 0.06 | 0.04 | 000067 |
| – | 2022-06-04 | R | 0.01 | 0.03 | 0.00 | 0.00 | 0.01 | 0.03 | 090157 |
| HD154892 | 2022-06-22 | Rx | 0.18 | 0.03 | 0.05 | 0.03 | 0.13 | 0.04 | NOT; 000157 |
| HD154892 | 2022-07-23 | Rx | 0.20 | 0.04 | 0.05 | 0.03 | 0.15 | 0.05 | NOT; 000157 |
| G191B2B | 2022-11-24 | R | 0.06 | 0.06 | 0.09 | 0.05 | -0.03 | 0.07 | NOT; 000157 |
| G191B2B | 2022-12-22 | R | 0.08 | 0.10 | 0.09 | 0.05 | -0.01 | 0.11 | NOT; 000157 |
| HD65629 | 2024-01-12 | IF443 | 0.31 | 0.04 | 0.01 | 0.00 | 0.31 | 0.04 | NOT; 000157 |
| HD65629 | 2024-01-12 | IF684 | 0.48 | 0.03 | 0.01 | 0.00 | 0.48 | 0.03 | NOT; 000157 |
| G191B2B | 2024-11-28 | R | 0.05 | 0.03 | 0.09 | 0.05 | -0.04 | 0.06 | NOT; 000157 |
| G191B2B | 2024-12-29 | Rx | 0.12 | 0.04 | 0.09 | 0.05 | 0.03 | 0.06 | NOT; 000157 |
| Star | Date | Filter | P, | P, | Pcat, | Pcat, | (P-Pcat), | (P-Pcat), | Notes |
| HD183143 | 2014-12-20 | IF890 | 4.56 | 0.06 | 4.72 | – | -0.17 | 0.06 | UKIRT; 000067 |
| HD183143 | 2014-12-20 | IF890 | 4.55 | 0.06 | 4.72 | – | -0.17 | 0.06 | UKIRT; 090157 |
| HD43384 | 2015-03-18 | IF890 | 2.30 | 0.02 | 2.22 | – | 0.08 | 0.02 | UKIRT; 000067 |
| HD183143 | 2015-03-21 | IF890 | 4.68 | 0.04 | 4.72 | – | -0.04 | 0.04 | UKIRT; 000067 |
| HD183143 | 2015-04-23 | IF890 | 4.56 | 0.06 | 4.72 | – | -0.17 | 0.06 | UKIRT; 000067 |
| HD183143 | 2015-04-23 | IF890 | 4.55 | 0.06 | 4.72 | – | -0.17 | 0.06 | UKIRT; 090157 |
| HD183143 | 2015-04-26 | IF890 | 4.56 | 0.06 | 4.72 | – | -0.17 | 0.06 | UKIRT; 000067 |
| HD183143 | 2015-04-26 | IF890 | 4.55 | 0.06 | 4.72 | – | -0.17 | 0.06 | UKIRT; 090157 |
| BDp59389 | 2017-12-15 | R | 6.47 | 0.02 | 6.43 | 0.02 | 0.04 | 0.03 | NOT; 000067 |
| BDp59389 | 2017-12-15 | R | 6.99 | 0.02 | 6.43 | 0.02 | 0.56 | 0.03 | NOT; 090157 |
| HD161056 | 2018-03-11 | R | 3.89 | 0.04 | 4.01 | 0.03 | -0.13 | 0.05 | NOT; 000067 |
| HD161056 | 2018-03-11 | R | 3.74 | 0.03 | 4.01 | 0.03 | -0.27 | 0.04 | NOT; 090157 |
| BDp59389 | 2019-10-25 | Rx | 6.34 | 0.04 | 6.43 | 0.02 | -0.09 | 0.05 | NOT; 000067 |
| BDp59389 | 2019-10-25 | Rx | 6.25 | 0.04 | 6.43 | 0.02 | -0.18 | 0.05 | NOT; 090157 |
| HD25443 | 2020-04-18 | Rx | 4.71 | 0.04 | 4.73 | 0.05 | -0.02 | 0.06 | NOT; 000067 |
| HD25443 | 2020-04-18 | Rx | 4.67 | 0.04 | 4.73 | 0.05 | -0.07 | 0.06 | NOT; 090157 |
| HD25443 | 2020-04-25 | Rx | 4.67 | 0.03 | 4.73 | 0.05 | -0.07 | 0.05 | NOT; 000067 |
| HD25443 | 2020-04-25 | Rx | 4.69 | 0.03 | 4.73 | 0.05 | -0.04 | 0.05 | NOT; 090157 |
| HD161056 | 2020-04-30 | Rx | 3.91 | 0.03 | 4.01 | 0.03 | -0.11 | 0.05 | NOT; 000067 |
| HD161056 | 2020-04-30 | Rx | 3.85 | 0.03 | 4.01 | 0.03 | -0.17 | 0.05 | NOT; 090157 |
| HD161056 | 2022-06-22 | Rx | 4.01 | 0.00 | 4.01 | 0.00 | -0.00 | 0.00 | NOT; 000157 |
| HD161056 | 2022-07-23 | Rx | 3.87 | 0.00 | 4.01 | 0.00 | -0.14 | 0.00 | NOT; 000157 |
| HD251204 | 2022-11-24 | R | 4.84 | 0.00 | 4.72 | 0.04 | 0.12 | 0.04 | NOT; 000157 |
| HD251204 | 2022-12-22 | R | 4.90 | 0.00 | 4.72 | 0.04 | 0.18 | 0.04 | NOT; 000157 |
| HD25443 | 2024-01-12 | IF443 | 5.49 | 0.00 | 5.23 | 0.09 | 0.26 | 0.09 | NOT; 000157 |
| HD25443 | 2024-01-12 | IF642 | 4.74 | 0.00 | 4.73 | 0.05 | 0.01 | 0.05 | NOT; 000157 |
| HD25443 | 2024-01-14 | IF642 | 4.77 | 0.00 | 4.73 | 0.05 | 0.04 | 0.05 | NOT; 000157 |
| HD204827 | 2024-04-10 | Rx | 4.76 | 0.00 | 4.89 | 0.03 | -0.13 | 0.03 | NOT; 000157 |
| HD161056 | 2024-04-10 | Rx | 3.92 | 0.00 | 4.01 | 0.03 | -0.10 | 0.03 | NOT; 000157 |
| Hiltner960 | 2024-04-10 | Rx | 4.97 | 0.00 | 5.21 | 0.03 | -0.24 | 0.03 | NOT; 000157 |
| VICyg12 | 2024-04-10 | Rx | 7.96 | 0.00 | 7.89 | 0.04 | 0.07 | 0.04 | NOT; 000157 |
| HD283812 | 2024-04-10 | Rx | 4.74 | 0.00 | 4.73 | 0.05 | 0.01 | 0.05 | NOT; 000157 |
| HD283812 | 2024-11-28 | R | 4.77 | 0.00 | 4.73 | 0.05 | 0.04 | 0.05 | NOT; 000157 |
Appendix B Polarised spectra of high polarisation standards.
In this appendix, we present the polarised spectra of the high polarisation standard stars: HD 7927, HD 25443, HD 161056, HD 19820, HD 204827, BD +25 727, HD 183143, BD +59 389, HD 154445, and HD 155197. The orange, red, and green dots represent the catalogue values taken from 1992AJ....104.1563S; 2021AJ....161...20P; 1983A&A...121..158S; 1990AJ.....99.1243T; 1982ApJ...262..732H, and the data presented on the Nordic Optical Telescope’s website888Source: https://www.not.iac.es/instruments/turpol/std/bd25727_note.html. The red curve represents Serkowski’s law, and the coefficients of Serkowski’s law are presented in Table LABEL:tab.Serkowski.
Appendix C 3D polarisation maps of high polarisation standards.
To demonstrate how the position angle of highly polarised standard stars is oriented, we created 3D maps. In Fig. 19, we present 3D polarisation maps in the vicinity of observed high-polarisation standards. We used data from 9286 Stars: An Agglomeration of Stellar polarisation Catalogues (2000AJ....119..923H). The colour of each star corresponds to its distance, based on data from Gaia Early Data Release 3 (2021AJ....161..147B). To match the data between the two catalogues, we used TOPCAT (2020ASPC..522...67T). The distance to BD +25 727 is unclear (presented with a blue dot in Fig. 19). High-polarisation standards are marked with a star symbol and are located at the centre of the images. The blue line of the high-polarisation standards represents catalogue values as presented in Table 1. In most cases, the position angle of the observed high-polarised standard star is aligned with the position angles of neighbouring stars. An exception to this is observed in HD 183143. The background image represents 100 m dust emission maps (1998ApJ...500..525S).